A car is moving with a constant velocity of 8 and then accelerates at a rate of 2 m/s/s for 3 seconds. Determine the displacement (in whole meters) of the car during its acceleration.
1
Expert's answer
2020-09-03T14:10:48-0400

The displacement of the car is given by the kinematic law:


"d = v_0t + \dfrac{at^2}{2}"

where "v_0 = 8m\/s" is the initial velocity, "a = 2m\/s^2" is the acceleration rate and "t = 3s" is time. Thus, obtain:


"d = 8\cdot 3 + \dfrac{2\cdot 3^2}{2} = 33 m"

Answer. 33 m.